After final preparations on both sides of the hatches (closed yesterday at 2:42pm EST), Discovery this morning undocked smoothly at 5:32am from PMA-2 (Pressurized Mating Adapter 2), after a total docked time of 10d 21h 52min. (For undocking, the station was turned ~180 deg to +XVV ZLV (+x-axis in velocity vector, z-axis in local vertical) at 4:28am, put briefly on free drift for the undocking, and then maneuvered to 10A Stage attitude of +XVV TEA attitude at 6:15am.)
KSC landing is nominally expected on 11/7 (Wednesday) at ~1:02pm EST. (If the landing occurs as planned, Discovery's mission duration will be 15d 2h 24m.)
Aboard the station, the crew is enjoying a light-duty day. Wakeup for CDR Peggy Whitson, FE-1 Yuri Malenchenko and FE-2 Dan Tani was at ~1:40am. Sleep time tonight: 4:30pm EST.
After the CDR prepared the auditory test equipment, she, FE-1 Malenchenko and FE-2 Tani took the periodic O-OHA (on-orbit hearing assessment) test, a 30-min. NASA environmental health systems examination to assess the efficacy of acoustic countermeasures, using a special MEC laptop application. It was the first session for the three crewmembers. (The O-OHA audiography test involves minimum audibility measurements for each ear over a wide range of frequencies (0.25-10 kHz) and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, Bose ANC headsets and the SLM (sound level meter). To conduct the testing, the experimenter is supported by special EarQ software on the MEC, featuring an up/down-arrow-operated slider for each test frequency that the crewmember moves to the lowest sound pressure level at which the tone can still be heard. The baseline test is required not later than about Flight Day 14 for each new Expedition and is then generally performed once per month. Note: There have been temporary hearing deficits documented on some U.S. and Russian crewmembers, all of which recovered to pre-mission levels.)
With the U.S. CDRA (Carbon Dioxide Removal Assembly) deactivated by the ground late last night (~10:30pm EST) and its cooling no longer required, Whitson demated and took down the ITCS LTL (Internal Thermal Control System/Low Temperature Loop) jumper at the CDRA-supporting LAB1D6 rack.
Both Dan Tani and Peggy Whitson used the DCS760 digital cameras and PD-100 camcorder to document the undocking, backing away and separation of the Discovery. Tani later conducted the playback for transmitting the footage to the ground.
FE-1 Malenchenko performed the periodic MLI (Multi-Layer Insulation) Nikon D2X photography on Soyuz TMA-11/15S, docked at the FGB nadir port, from the DC1 EVA hatch 2 window, to assess the condition of the spacecraft's MLI (Multi-Layer Insulation) material.
After the crew reconfigured the PMA-2-to-Lab vestibule last night in preparation for PMA-2 demate & relocation to Node-2, removed the CBM (Common Berthing Mechanism) ground strap and connected the CBM bolt actuator to load cell sensor, Yuri opened the Lab hatch to the PMA-2 today, to verify that all Kapton tape is removed from the PMA-2 grille and that the Velcro grille cover is off of the PMA-2 hard duct.
CDR Whitson later performed the remaining preparatory PMA-2 tasks. (Work items consisted of disconnecting the O 2 & N 2 recharge lines, installing the four CBM CPAs (Controller Panel Assemblies (CPA), installing the CBCS (Centerline Berthing Camera System) target assembly, removing all stowage & tools from PMA-2, verifying CBM hardware config, conducting photo documentation, and installing the Lab forward hatch thermal cover.)
Afterwards, Dan Tani removed the PMA-2 ventilation air duct, closed the Lab forward hatch, depressurized the PMA-2 and performed a leak check of the Lab fwd hatch (several times).
The new Exp-16 crew, Peggy, Yuri & Dan, performed a one-hour OBT (onboard training) emergency egress drill. (Purpose of the refresher drill was to familiarize the entire crew with Node-2 hardware and valves used in emergency situations, familiarize Dan Tani with the locations of ISS hardware and the positions of valves used in emergency situations as well as the translation routes to the Soyuz crew return vehicle, working through the Russian Segment (RS) hardware deactivation procedures, and practicing crew interactions in emergency situations.)
In the Airlock, the two spacewalkers on EVA-5 next Friday (11/9), Whitson and Malenchenko, relocated and configured their EMU (Extravehicular Mobility Unit) spacesuits, i.e., EMU #3006 on the forward EDDA (EMU Don Doff Assembly), #3018 on the aft EDDA. (After terminating recharge of the suit batteries and regeneration of the METOX (Metal Oxide) filter canisters, Peggy and Yuri installed both batteries and canisters in the EMUs. Afterwards, they verified proper configuration for helmet lights, video cameras (ERCAs), identification stripes, wrist mirrors and cuff checklists.)
The CDR also deconfigured the BPSMU (Battery Powered Speaker Microphone Unit) and its long drag-through cable, used during the docked phase, and stowed the equipment.
Malenchenko completed the post-undocking reconfiguration of the Russian telephone/telegraph subsystem (STTS) to its pre-docked settings, from its primary string back to nominal mode on the backup string. This also severed the VHS (UHF) channel to the receding Shuttle Orbiter and restored the RSA-2 S/G (Space-to-Ground) comm configuration on Panel 3. (The ""Voskhod-M"" STTS enables telephone communications between the SM, FGB, DC1 Docking Compartment and U.S. segment (USOS), and also with users on the ground over VHF channels selected by an operator at an SM comm panel, via STTS antennas on the SM's outside. There are six comm panels in the SM with pushbuttons for accessing any of three audio channels, plus an intercom channel. Other modes of the STTS include telegraphy (teletype), EVA voice, emergency alarms, Packet/Email, and TORU docking support.)
FE-2 Tani rerouted an EWIS (External Wireless Instrumentation System) 1553 cable in the Lab, disconnecting it, then reconnecting it behind another panel and closing the panel cover.
As is regular procedure for newly arrived cosmonauts, the FE-1 spent some time filling out the questionnaire for the standard Russian biomedical Braslet-M/Anketa (""bracelet/questionnaire"") test procedure. (The objective is for Yuri to evaluate a number of ""bracelet"" cuffs for their usefulness in suppressing the adverse effects of micro-G for the ""newcomer"" aboard the station during the acute phase of adaptation to weightlessness, if there are such indications. The ""bracelets"" are compression cuffs attached to a belt and worn on the upper thighs over the coveralls, intended as countermeasures against the initial micro-G effects of blood filling (vascularity) in the upper torso (heaviness and blood pulsation in the head), facial puffiness, nasal stuffiness, painful eye movement, and vestibular disorders (dizziness, nausea, vomiting). They create artificial blood accumulation in the upper thirds of the thighs, causing some of the circulating blood volume to relocate from the upper body to the lower extremities, thereby (hopefully) correcting the adverse hemodynamic effect of micro-G and thus improving the crewmember's working capability. The actual compression cuff in the Braslet units is a combination of alternating multi-layer tensile and non-tensile elements, whose distension by body movements creates elastic forces that produce the necessary pressure on the body surface.)
At ~8:50am, Yuri supported the ground's reactivation of the Elektron O 2 generator at 32 amps by monitoring the external temperature of its secondary purification unit (BD) for the first 10 minutes of operations to ensure that there was no overheating. (During nominal operations a gas analyzer is utilized to detect hydrogen (H 2) in the O 2 line (which could cause overheating) but is not included in the control algorithm until 10 minutes after Elektron startup.)
Later, the FE-1 performed the routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM (Service Module), including routine ASU toilet facilities replacings.
ISS crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR, FE-2), TVIS treadmill (FE-1, FE-2), RED resistive exercise device (CDR), and VELO bike with bungee cord load trainer (FE-1).
Yuri then transferred the crew's exercise data file to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).
Payload Transfers: As part of ISS-to-Shuttle transfers yesterday, the crew packed science samples and ICEPACs in two DCBs (Double Coldbags) on Discovery before hatch closure. When Clay Anderson ran into some difficulties packing one of the bags during these activities, he was instructed to remove seven NUTRITION samples from one of the bags and re-stow them in the MELFI (Minus Eighty-Degree Laboratory Freezer for ISS) until the next flight.
Water Transfers: A total of 10 CWCs (Contingency Water Containers) were transferred from Discovery to the station. CDR Whitson reported that a CWC (#1006), previously filled and transferred, was found to be leaking. The CWC was returned to the Shuttle, to be vented overboard during a post-undock water dump.
No CEO (Crew Earth Observations) target uplinked for today.
ISS Orbit (as of this morning, 8:55am EST (= epoch)):
Mean altitude -- 342.2 km
Apogee height -- 344.1 km
Perigee height -- 340.3 km
Period -- 91.38 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0002835
Solar Beta Angle -- -52.7 deg (magnitude decreasing)
Orbits per 24-hr. day -- 15.76
Mean altitude gain in the last 24 hours -- 89 m (Shuttle-caused delta-V)
Revolutions since FGB/Zarya launch (Nov. 98) -- 51308